OpenStudy (anonymous):

Given the functions m(x) = 4x − 11 and n(x) = x − 10, solve m[n(x)] and select the correct answer below.

OpenStudy (anonymous):

@ospreytriple

OpenStudy (anonymous):

m[n(x)] = 4x − 51 m[n(x)] = 4x − 29 m[n(x)] = 4x2− 51 m[n(x)] = 4x2 − 29

OpenStudy (anonymous):

OK. Slightly trickier, but the method is exactly the same.

OpenStudy (anonymous):

If we want m(n(x)), then you must take n(x) first. That is given as \[n(x) = x - 10\]Now, to find \(m(n(x))\), you must calculate \(m(x-10)\). Using the function m(x), substitute \(x-10\) for \(x\).

OpenStudy (anonymous):

In other words\[m \left( n \left( x \right) \right) = m \left( x-10 \right) = 4\left( x-10 \right)-11\]What do you get?

OpenStudy (anonymous):

ummm what are putting for x

OpenStudy (anonymous):

Nothing, just leave it as x. Notice there are x's in all the answers.

OpenStudy (anonymous):

irdk im sry

OpenStudy (anonymous):

No problem. So you need to figure out 4(x-10) - 11. Take it one bit at a time. Can you multiply 4(x-10)?

OpenStudy (anonymous):

4x-40

OpenStudy (anonymous):

distribute

OpenStudy (anonymous):

Perfect. Now subtract 11 from that answer, i.e. 4x - 40 - 11.

OpenStudy (anonymous):

Combine the -40 and the -11.

OpenStudy (anonymous):

sry it is --51 @ospreytriple

OpenStudy (anonymous):

Right. So your answer is 4x - 51. OK?

OpenStudy (anonymous):

yes sry i was doing something with my teacher over the phone @ospreytriple

OpenStudy (anonymous):

I'm in flvs (florida virtual school)